# Statistical properties of Markov shifts (part I)

### v1

@@ Abstract.
Our results apply to Markov shifts in random dynamical environment, products of random non-stationary positive matrices and other operators, random Lyapunov exponents, non-autonomous non-uniformly expanding transformations, as well as several processes arising in statistics and applied probability like linear processes, inhomogeneous iterated random functions and GARCH processes. Most of these examples seem to only be treated in literature for iid $X_{j}$ and here we are able to drop both the stationarity and the independence assumptions. However, even in the classical setup of Anosov maps or subshifts of finite type our results seem to be new when working with measures of maximal entropy since we can consider observables which are only Hölder continuous on average.
@@ 1. Introduction
Let $(Y_{j})$ be an independent sequence of zero mean square integrable random variables, and let $S_{n}=\sum_{j=1}^{n}Y_{j}$. The classical CLT states that if $\lim_{n\to\infty}\|S_{n}\|_{L^{2}}=\infty$ then $S_{n}/\|S_{n}\|_{L^{2}}$ converges in distribution to the standard normal law if and only if the Lindeberg condition [^1] Namely that $\lim_{n\to\infty}{\sigma}_{n}^{-2}\sum_{j=1}^{n}{\mathbb{E}}[Y_{j}^{2}{\mathbb{I}}(|Y_{j}|\geq\varepsilon{\sigma}_{n})]=0$ for all $\varepsilon>0$. holds. In particular, when $Y_{j}$ are identically distributed and non-constant then the weak limit is Gaussian. Note that when ${\sigma}_{n}\not\to\infty$ then by Kolmogorov’s three-series theorem $S_{n}$ converges almost surely (and also in $L^{2}$) and so there is no weak limit in general.
@@ 1.1. CLT rates for independent summands
In this paper we are interested in the behavior of non-stationary processes. In the context of independent summands this means that $Y_{j}$ are not identically distributed. In that case the classical result of Berry and Esseen shows that $\Delta_{n}$ is at most of order ${\sigma}_{n}^{-3}\sum_{j=1}^{n}{\mathbb{E}}[|Y_{j}|^{3}]$. Similarly to the iid case, when dropping the requirement of identical distribution when $\sigma_{n}\to\infty$ then $\Delta_{n}=o({\sigma}_{n}^{-1})$ if and only if $(Y_{j})$ is not reducible to a lattice valued sequence and $\sum_{j=1}^{n}{\mathbb{E}}[Y_{j}^{3}]=o({\sigma}_{n}^{2})$ (see [23]). All of the above results are obtained by using Fourier analysis methods applied with the Fourier transform of the measure induced by $F_{n}$ (i.e. the characteristic function of $S_{n}/{\sigma}_{n}$).
@@ 1.2.1. Stationary sequences
Note that the CLT itself is due to Dobrushin [22], see the next section. Since then optimal CLT rates $O(n^{-1/2})$ (aka Berrry Esseen theorems) were obtained for many classes of stationary processes $Y_{k}$ under some mixing (weak dependence) and moment assumptions on $Y_{k}$, see [46, 37, 36, 35, 61, 63] for a few general approaches for chaotic dynamical systems, Markov chains, Bernoulli shifts and bounded $\phi$-mixing sequences. Such result have applications in other areas of probability and statistics like products of random matrices (see [18]), iterated function systems and other processes arising in statistics and applied probability [19, 35], and many others. Of course, there are many other results in literature but it is beyond the scope of this paper to provide a full list. In the stationary setting our results seem to be new as the setting of functions that depend on the entire path of a Markov chain was not treated, but for sufficiently regular functions we expect such results to follow from [46]. However, note that for Markov measures (including measures of maximal entropy (MME)) on subshifts of finite type or Anosov maps (via symbolic representations) our results apply to functions $f$ which are only Hölder continuous on average. This was not treated in literature and does not seem to immediately follow from existing results.
@@ 1.3.2. Random dynamical systems
Here we focus our attention on our applications to Markov shifts in random dynamical environment (see Section 2.5.2). One can view this setup as a special case of a random dynamical system (RDS). RDS are motivated by real life phenomenon of random noise which make a given system non-stationary in nature. Ergodic theory of RDS has attracted a lot of attention in the past decades, see [2, 13, 16, 48, 53, 49]. We refer to the introduction of [49, Chapter 5] for a historical discussion and applications to, for instance, statistical physics, economy and meteorology etc. The literature on statistical properties (i.e. limit theorems) of random dynamical systems exploded in recent years. Let us mention only a few results which are most relevant to our work. In [12] central limit theorems were studied for Markov chains in random dynamical environment (as opposed to Markov shifts). In [51] central limit theorems were studied for a variety of random dynamical systems, while in [50] large deviations were obtained. In [27, 28, 29] central limit theorems, large deviations and almost sure invariance principle were obtained. Berry-Esseen theorems were obtained in [31, 38, 41]. In the past two decades the number of papers on the asymptotic behavior of random dynamical systems has exploded, and so again we will not make an attempt to even provide a full list. Our contribution to the theory of random dynamical systems is that we can consider functionals which depend on the entire path of the Markov chain (in the random environment), which includes applications to many other natural processes in random environment.
@@ 2.1. Mixing and approximation conditions
Let $(X_{j})_{j\in{\mathbb{Z}}}$ be a Markov chain defined on some probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$. For all $\infty\leq k\leq\ell\leq\infty$ let us denote by ${\mathcal{F}}_{k,\ell}$ the $\sigma$}-algebra generated by $X_{s}$ for all finite $k\leq s\leq\ell$. Let $1\leq q\leq p\leq\infty$, and recall that (see [8, Ch. 4]) the (reverse) $\varpi_{q,p}$ mixing (weak dependence) coefficients associate with the chain are given by
@@
Recall that (see [8, Ch.4]), the more familiar $\rho,\phi$ and $\psi$ mixing coefficients can be written as
@@ 2.2. Moments and mixing type assumptions needed for optimal CLT rates
Let $p,q\geq 1$ and $s\geq 3$ be such that $\frac{1}{p}=\frac{1}{s}+\frac{1}{q}$ (so $q>p$). Suppose that (2.3) holds with these $q$ and $p$. Let $a>s$. Suppose that there exist $\delta\in(0,1)$ and $c_{0},\zeta>0$ such that $\|f_{j}\|_{j,a,s,\delta}\leq c_{0}(j+1)^{\zeta},j\geq 0$. Moreover, there exist $c_{1},n_{0}>0$ and $\varepsilon>0$ such that ${\sigma}_{n}^{2}\geq c_{1}n^{(1+\zeta)\frac{6}{a-3}+2\zeta+\varepsilon}$ for all $n\geq n_{0}$, where when $a=\infty$ we set $\frac{6}{a-3}=0$. Under this assumption we set $k=\infty$.
@@
$\varepsilon_{0}=\sup\left\{\varepsilon>0:\,\liminf_{n\to\infty}\frac{{\sigma}_{n}^{2}}{n^{\varepsilon+2\zeta+\frac{6(1+\zeta)}{a-3}}}>0\right\}.$
@@
Let $p,q\geq 2$ and $s\geq 2$ be such that $\frac{1}{p}=\frac{1}{s}+\frac{1}{q}$ (so $q>p$). Let us assume that there exists $a<p/2$ such that $1/3=1/p+1/a$. Suppose that (2.3) holds with these $q$ and $p$ and there exists $\delta\in(0,1)$ such that $\sup_{j}\|f_{j}\|_{j,p,s,\delta}<\infty$. Moreover, ${\sigma}_{n}\to\infty$ and for all $n$ and $A$ large enough there exists a variance partition corresponding to $A$ such that $\sup_{\ell}\max_{B\subset B_{\ell}}\|S_{B}\|_{L^{p}}<\infty$ (where $B$ is a sub-interval whose left end point is the same as $B_{\ell}$). Under this assumption we set $k=\infty$.
@@ 2.2.1. A moment condition for block decompositions
We have $\lim_{n\to\infty}{\sigma}_{n}=\infty$. Moreover, either $V:=\liminf_{n\to\infty}\frac{1}{n}\text{Var}(S_{n})>0$ or for all $n$ and $A>1$ large enough there exists a variance partition $(B_{\ell,n})$ of $\{0,1,...,n-1\}$ such that
(2.9) $\tilde{L}_{3,n}:=\sum_{j=1}^{k_{n}}{\mathbb{E}}\left[|S_{B_{j}}-{\mathbb{E}}[S_{B_{j,n}}]|^{3}\right]=O({\sigma}_{n}^{2}).$
Assumption 2.12 is very similar in spirit to (2.8), except that we need to consider the sums of third moments along blocks $B_{j}$. The reason is that our methods are based on a block partition argument which is crucial in overcoming the fact that ${\sigma}_{n}^{2}$ can grow sublinearly fast.
@@
(1) if (2.2) holds with $p=\infty$ and some $1\leq q\leq p$ and $\sup_{j}\|f_{j}\|_{j,\infty,\infty,\delta}<\infty$ for some $\delta\in(0,1)$.
(2) $f_{j}$ is a reversed martingale difference, (2.2) holds with some $1\leq q\leq p$, $\sup_{j}\|f_{j}^{2}\|_{j,q,p,\delta}<\infty$, and there exists a constant $C>0$ such that
@@
(3) $f_{j}$ is a forward martingale difference, (2.2) holds with some $1\leq p\leq q$, $\sup_{j}\|f_{j}^{2}\|_{j,p,q,\delta}<\infty$, and (2.10) holds with $u$ being the conjugate exponent of $q$.
(4) $\mu_{j}(f_{j})=0$ for all $j$, and $\varpi_{p,q}(n)\to 0$ for some conjugate exponents $q,p$ with $p\leq 2$ and for some $\delta\in(0,1)$ we have
@@ 2.3. The growth of the variance, Livsic regularity and the CLT
Next we address the CLT. First, we describe what is essentially known in literature in our setup. The following result follows by the discussion in [42, Section 7.2].
@@
(2.14) $\lim_{n\to\infty}{\sigma}_{n}^{-2}\sum_{j=0}^{n-1}{\mathbb{E}}[|f_{j}|^{2}{\mathbb{I}}(|f_{j}|\geq{\varepsilon}{\sigma}_{n})]=0.$
@@ 2.4. Optimal CLT rates and moment estimates
Let one of Assumptions 2.6, 2.7, 2.10 or 2.11 be in force and let $k$ be as described at the assumptions. Under one of Assumption 2.6 and 2.11 let $u=0$. While under Assumption 2.7 let $u=1-\varepsilon(1-\varepsilon)/2$ and under Assumption 2.10 let $u=2a/p<1$. Suppose also that Assumption 2.12 holds.
(i) for all finite $0\leq s\leq k-1$ there is a constant $C_{s}$ such that
@@
(ii) for all $q>\frac{1}{k}$ we have $\left\|F_{n}-\Phi\right\|_{L^{q}(dx)}=O({\sigma}_{n}^{-(1-u)}).$
(iii) for all finite $1\leq s\leq k-1$ there is a constant $C_{s}$ such that for every absolutely continuous function $h:{\mathbb{R}}\to{\mathbb{R}}$ such that $H_{s}(h):=\int\frac{|h^{\prime}(x)|}{1+|x|^{s}}dx<\infty$ we have
@@
Let one of Assumptions 2.6, 2.7, 2.10 2.11 be in force and let $k$ be as described at the assumptions. Suppose also that Assumption 2.12 holds. Then, for every finite b<k1b<k-1 we have
@@
Using the main results in [30] and a block partition argument we can get rates of order $O(\sigma_{n}^{1/2+\varepsilon})$ in the almost sure invariance principle (ASIP), that is we can couple $S_{n}f$ with a Brownian motion $B(t)$ such that $|S_{n}f-{\mathbb{E}}[S_{n}f]-B(\sigma_{n}^{2})|=O(\sigma_{n}^{1/2+\varepsilon})$, almost surely. This implies the functional central limit theorem and the law of iterated logarithm, for example (see [60]). It also implies other limit laws like the Arcsine Law and the law of records, see [11, Appendix C]. However, the rates $O(\sigma_{n}^{1/2+\varepsilon})$ are suboptimal compared with the rates $O(\ln n)$ established recently for various stationary processes (see [17]). In our setup the optimal rates should be $O(\ln\sigma_{n})$, which seems to require a different approach. Because of these reasons we decided only to remark on the above ASIP rates, and to address the problem of getting optimal rates elsewhere.

### v2

@@ Abstract.
Our results apply to Markov shifts in random dynamical environment, products of random non-stationary positive matrices and other operators, random Lyapunov exponents, non-autonomous non-uniformly expanding transformations, as well as several processes arising in statistics and applied probability like linear processes, inhomogeneous iterated random functions and GARCH processes. Most of these examples seem to only be treated in literature for iid $X_{j}$ and here we are able to drop both the stationarity and the independence assumptions. However, even in the classical setup of Anosov maps, subshifts of finite type and Gibbs-Markov maps our results seem to be new when working with measures of maximal entropy since we can consider observables which are only Hölder continuous on average.
@@ 1. Introduction
Let $(Y_{j})$ be an independent sequence of zero mean square integrable random variables, and let $S_{n}=\sum_{j=1}^{n}Y_{j}$. The classical CLT states that if $\lim_{n\to\infty}\sigma_{n}=\infty,\sigma_{n}=\|S_{n}\|_{L^{2}}$ then $S_{n}/\sigma_{n}$ converges in distribution to the standard normal law if and only if the Lindeberg condition [^1] Namely that $\lim_{n\to\infty}{\sigma}_{n}^{-2}\sum_{j=1}^{n}{\mathbb{E}}[Y_{j}^{2}{\mathbb{I}}(|Y_{j}|\geq\varepsilon{\sigma}_{n})]=0$ for all $\varepsilon>0$. holds. In particular, when $Y_{j}$ are identically distributed and non-constant then the weak limit is Gaussian. Note that when ${\sigma}_{n}\not\to\infty$ then by Kolmogorov’s three-series theorem $S_{n}$ converges almost surely (and also in $L^{2}$) and so there is no weak limit in general.
@@ 1.1. CLT rates for independent summands
In this paper we are interested in the behavior of non-stationary processes. In the context of independent summands this means that $Y_{j}$ are not identically distributed. In that case the classical result of Berry and Esseen shows that $\Delta_{n}$ is at most of order ${\sigma}_{n}^{-3}\sum_{j=1}^{n}{\mathbb{E}}[|Y_{j}|^{3}]$. Similarly to the iid case, when dropping the requirement of identical distribution when $\sigma_{n}\to\infty$ then $\Delta_{n}=o({\sigma}_{n}^{-1})$ if and only if $(Y_{j})$ is not reducible to a lattice valued sequence and $\sum_{j=1}^{n}{\mathbb{E}}[Y_{j}^{3}]=o({\sigma}_{n}^{2})$ (see [24]). All of the above results are obtained by using Fourier analysis methods applied with the Fourier transform of the measure induced by the law of $S_{n}/{\sigma}_{n}$ (i.e. the characteristic function of $S_{n}/{\sigma}_{n}$).
@@ 1.2.1. Stationary sequences
Note that the CLT itself is due to Dobrushin [23], see the next section. Since then optimal CLT rates $O(n^{-1/2})$ (aka Berrry Esseen theorems) were obtained for many classes of stationary processes $Y_{k}$ under some mixing (weak dependence) and moment assumptions on $Y_{k}$, see [47, 38, 37, 36, 62, 64] for a few general approaches for chaotic dynamical systems, Markov chains, Bernoulli shifts and bounded $\phi$-mixing sequences. Such result have applications in other areas of probability and statistics like products of random matrices (see [19]), iterated function systems and other processes arising in statistics and applied probability [20, 36], and many others. Of course, there are many other results in literature but it is beyond the scope of this paper to provide a full list. In the stationary setting our results seem to be new as the setting of functions that depend on the entire path of a Markov chain was not treated, but for sufficiently regular functions we expect such results to follow from [47]. However, note that for Markov measures (including measures of maximal entropy (MME)) on subshifts of finite type, Gibbs Markov maps or Anosov maps (via symbolic representations) our results apply to functions $f$ which are only Hölder continuous on average. This was not treated in literature and does not seem to immediately follow from existing results.
@@ 1.3.2. Random dynamical systems
Here we focus our attention on our applications to Markov shifts in random dynamical environment (see Section 2.5.2). One can view this setup as a special case of a random dynamical system (RDS). RDS are motivated by real life phenomenon of random noise which make a given system non-stationary in nature. Ergodic theory of RDS has attracted a lot of attention in the past decades, see [3, 14, 17, 49, 54, 50]. We refer to the introduction of [50, Chapter 5] for a historical discussion and applications to, for instance, statistical physics, economy and meteorology etc. The literature on statistical properties (i.e. limit theorems) of random dynamical systems exploded in recent years. Let us mention only a few results which are most relevant to our work. In [13] central limit theorems were studied for Markov chains in random dynamical environment (as opposed to Markov shifts). In [52] central limit theorems were studied for a variety of random dynamical systems, while in [51] large deviations were obtained. In [28, 29, 30] central limit theorems, large deviations and almost sure invariance principle were obtained. Berry-Esseen theorems were obtained in [32, 39, 42]. In the past two decades the number of papers on the asymptotic behavior of random dynamical systems has exploded, and so again we will not make an attempt to even provide a full list. Our contribution to the theory of random dynamical systems is that we can consider functionals which depend on the entire path of the Markov chain (in the random environment), which includes applications to many other natural processes in random environment.
@@ 2.1. Mixing and approximation conditions
Let $(X_{j})_{j\in{\mathbb{Z}}}$ be a Markov chain defined on some probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$. For all $\infty\leq k\leq\ell\leq\infty$ let us denote by ${\mathcal{F}}_{k,\ell}$ the $\sigma$}-algebra generated by $X_{s}$ for all finite $k\leq s\leq\ell$. Let $1\leq q\leq p\leq\infty$, and recall that (see [9, Ch. 4]) the (reverse) $\varpi_{q,p}$ mixing (weak dependence) coefficients associate with the chain are given by
@@
Recall that (see [9, Ch.4]), the more familiar $\rho,\phi$ and $\psi$ mixing coefficients can be written as
@@ 2.2. Moments and mixing type assumptions needed for optimal CLT rates
Let $p,q\geq 1$ and $s\geq 3$ be such that $\frac{1}{p}=\frac{1}{s}+\frac{1}{q}$ (so $q>p$). Suppose that (2.3) holds with these $q$ and $p$. Let $a>s$. Suppose that there exist $\delta\in(0,1)$ and $c_{0},\zeta>0$, $a>k_{0}\geq 3$ such that $\|f_{j}\|_{j,a,s,\delta}\leq c_{0}(j+1)^{\zeta},j\geq 0$. Moreover, there exist $c_{1},n_{0}>0$ and $\varepsilon>0$ such that ${\sigma}_{n}^{2}\geq c_{1}n^{(1+\zeta)\frac{2k_{0}}{a-k_{0}}+2\zeta+\varepsilon}$ for all $n\geq n_{0}$, where when $a=\infty$ we set $\frac{2k_{0}}{k_{0}-3}=0$. Under this assumption we set $k=k_{0}$.
@@
$\varepsilon_{0}=\sup\left\{\varepsilon>0:\,\liminf_{n\to\infty}\frac{{\sigma}_{n}^{2}}{n^{\varepsilon+2\zeta+\frac{2k_{0}(1+\zeta)}{a-k_{0}}}}>0\right\}.$
@@
Let $p,q\geq 2$ and $s\geq 2$ be such that $\frac{1}{p}=\frac{1}{s}+\frac{1}{q}$ (so $q>p$). Let us assume that there exists $3\leq k_{0}<a<p/2$ such that $1/k_{0}=1/p+1/a$. Suppose that (2.3) holds with these $q$ and $p$ and there exists $\delta\in(0,1)$ such that $\sup_{j}\|f_{j}\|_{j,p,s,\delta}<\infty$. Moreover, ${\sigma}_{n}\to\infty$ and for all $n$ and $A$ large enough there exists a variance partition corresponding to $A$ such that $\sup_{\ell}\max_{B\subset B_{\ell}}\|S_{B}\|_{L^{p}}<\infty$ (where $B$ is a sub-interval whose left end point is the same as $B_{\ell}$). Under this assumption we set $k=k_{0}$.
@@ 2.2.1. A moment condition for block decompositions
We have $\lim_{n\to\infty}{\sigma}_{n}=\infty$. Let $k$ be like in one of Assumptions 2.6, 2.7, 2.10 or 2.11, depending on the case. Let $k$ be like in one of Assumptions 2.6, 2.7, 2.10 or 2.11. Suppose that either $V:=\liminf_{n\to\infty}\frac{1}{n}\text{Var}(S_{n})>0$ or that for some finite $3\leq k_{0}\leq k$ for all $n$ and $A>1$ large enough there exists a variance partition $(B_{\ell,n})$ of $\{0,1,...,n-1\}$ such that
(2.9) $\tilde{L}_{k_{0},n}:=\sum_{j=1}^{k_{n}}{\mathbb{E}}\left[|S_{B_{j,n}f}-{\mathbb{E}}[S_{B_{j,n}f}]|^{k_{0}}\right]=O({\sigma}_{n}^{2}).$
When $k_{0}=3$ Assumption 2.12 is very similar in spirit to (2.8), except that we need to consider the sums of third moments along blocks $B_{j}$. The reason is that our methods are based on a block partition argument which is crucial in overcoming the fact that ${\sigma}_{n}^{2}$ can grow sublinearly fast.
@@
(1) if (2.2) holds with $p=\infty$ and some $1\leq q\leq p$ and $\sup_{j}\|f_{j}\|_{j,\infty,\infty,\delta}<\infty$ for some $\delta\in(0,1)$ the Assumption 2.12 holds with every finite $k_{0}$.
(2) Assumption 2.12 holds with $k_{0}=4$ if $f_{j}$ is a reversed martingale difference, (2.2) holds with some $1\leq q\leq p$, $\sup_{j}\|f_{j}^{2}\|_{j,q,p,\delta}<\infty$, and there exists a constant $C>0$ such that
@@
(3) Assumption 2.12 holds with $k_{0}=4$ if $f_{j}$ is a forward martingale difference, (2.2) holds with some $1\leq p\leq q$, $\sup_{j}\|f_{j}^{2}\|_{j,p,q,\delta}<\infty$, and (2.10) holds with $u$ being the conjugate exponent of $q$.
(4) Assumption 2.12 holds with $k_{0}=4$ if $\mu_{j}(f_{j})=0$ for all $j$, and $\varpi_{p,q}(n)\to 0$ for some conjugate exponents $q,p$ with $p\leq 2$ and for some $\delta\in(0,1)$ we have
@@ 2.3. The growth of the variance, Livsic regularity and the CLT
Next we address the CLT. First, we describe what is essentially known in literature in our setup. The following result follows by the discussion in [43, Section 7.2].
@@
(2.14) $\lim_{n\to\infty}{\sigma}_{n}^{-2}\sum_{j=0}^{n-1}{\mathbb{E}}[|f_{j}|^{2}{\mathbb{I}}(|f_{j}|\geq{\varepsilon}{\sigma}_{n})]=0$
@@ 2.4. Optimal CLT rates and moment estimates
Let one of Assumptions 2.6, 2.7, 2.10 or 2.11 be in force and let $k,k_{0}$ be as described in the assumptions. Under one of Assumption 2.6 and 2.11 let $u=0$. While under Assumption 2.7 let $u=1-\varepsilon(1-\varepsilon)/2$ and under Assumption 2.10 let $u=2a/p<1$. Suppose also that Assumption 2.12 holds.
(i) for all finite $0\leq s\leq k_{0}-1$ there is a constant $C_{s}$ such that
@@
(ii) for all $q>\frac{1}{k_{0}}$ we have $\left\|F_{n}-\Phi\right\|_{L^{q}(dx)}=O({\sigma}_{n}^{-(1-u)}).$
(iii) for all finite $1\leq s\leq k_{0}-1$ there is a constant $C_{s}$ such that for every absolutely continuous function $h:{\mathbb{R}}\to{\mathbb{R}}$ such that $H_{s}(h):=\int\frac{|h^{\prime}(x)|}{1+|x|^{s}}dx<\infty$ we have
@@
Let one of Assumptions 2.6, 2.7, 2.10 2.11 be in force and let $k,k_{0}$ be as described at the assumptions. Suppose also that Assumption 2.12 holds. Then, for every finite $b<k_{0}-1$ we have
@@
Using the main results in [31] and a block partition argument we can get rates of order $O(\sigma_{n}^{1/2+\varepsilon})$ in the almost sure invariance principle (ASIP), that is we can couple $S_{n}f$ with a Brownian motion $B(t)$ such that $|S_{n}f-{\mathbb{E}}[S_{n}f]-B(\sigma_{n}^{2})|=O(\sigma_{n}^{1/2+\varepsilon})$, almost surely. This implies the functional central limit theorem and the law of iterated logarithm, for example (see [61]). It also implies other limit laws like the Arcsine Law and the law of records, see [12, Appendix C]. However, the rates $O(\sigma_{n}^{1/2+\varepsilon})$ are suboptimal compared with the rates $O(\ln n)$ established recently for various stationary processes (see [18]). In our setup the optimal rates should be $O(\ln\sigma_{n})$, which seems to require a different approach. Because of these reasons we decided only to remark on the above ASIP rates, and to address the problem of getting optimal rates elsewhere.